# Dynamic Graph Transformer with Correlated Spatial-Temporal Positional Encoding

### v1

@@ Abstract.
Learning effective representations for Continuous-Time Dynamic Graphs (CTDGs) has garnered significant research interest, largely due to its powerful capabilities in modeling complex interactions between nodes. A fundamental and crucial requirement for representation learning in CTDGs is the appropriate estimation and preservation of proximity. However, due to the sparse and evolving characteristics of CTDGs, the spatial-temporal properties inherent in high-order proximity remain largely unexplored. Despite its importance, this property presents significant challenges due to the computationally intensive nature of personalized interaction intensity estimation and the dynamic attributes of CTDGs. To this end, we propose a novel Correlated Spatial-Temporal Positional encoding that incorporates a parameter-free personalized interaction intensity estimation under the weak assumption of the Poisson Point Process. Building on this, we introduce the Dynamic Graph Transformer with Correlated Spatial-Temporal Positional Encoding (CorDGT), which efficiently retains the evolving spatial-temporal high-order proximity for effective node representation learning in CTDGs. Extensive experiments on seven small and two large-scale datasets demonstrate the superior performance and scalability of the proposed CorDGT.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Knowledge representation and reasoningccs: Theory of computation Dynamic graph algorithms
@@ 1. Introduction
Graph Neural Networks (GNNs) (Veličković et al., 2018; Hamilton et al., 2017; Xu et al., 2019; Hou et al., 2022) have become a potent tool for analyzing diverse graph structures due to their effectiveness in learning low-dimensional graph representations. While early GNN research focused on static graphs, many real-world network data, such as the World Wide Web and social networks, exhibit evolving graph structures. The increasing prevalence of dynamic graph data has spurred researchers to adapt GNNs to handle dynamic graphs. Dynamic graphs can be categorized into Discrete-Time Dynamic Graphs (DTDGs) and Continuous-Time Dynamic Graphs (CTDGs). Recently, CTDGs have garnered more research attention compared to DTDGs, attributed to their flexible generalization capabilities and proficiency in modeling intricate and dynamic node interactions. To learn effective representations for CTDGs, researchers have explored various techniques, including temporal message passing (Xu et al., 2020b; Rossi et al., 2020; Kumar et al., 2019; Trivedi et al., 2019) and the incorporation of temporal intervals into random walks (Nguyen et al., 2018; Wang et al., 2021b; Jin et al., 2022), etc.
@@ 5.3. Ablation Studies
1. Set $\alpha=0$ in Eq. (7) 98.09$\pm$0.0 98.74$\pm$0.1 98.17$\pm$0.0 98.28$\pm$0.3 89.10$\pm$0.6 92.07$\pm$0.4 91.30$\pm$0.3 91.52$\pm$0.3
2. Set $\beta=0$ in Eq. (7) 98.10$\pm$0.0 98.69$\pm$0.1 97.87$\pm$0.0 98.18$\pm$0.2 89.55$\pm$0.6 90.80$\pm$1.7 90.46$\pm$0.3 94.28$\pm$0.8
3. Remove temporal distance 97.91$\pm$0.1 98.83$\pm$0.3 97.81$\pm$0.1 97.11$\pm$0.1 88.43$\pm$0.1 86.68$\pm$0.4 92.19$\pm$0.1 91.20$\pm$0.3
4. Remove spatial distance 97.81$\pm$0.0 98.52$\pm$0.1 97.62$\pm$0.0 97.36$\pm$0.4 92.47$\pm$0.1 91.43$\pm$0.5 90.15$\pm$0.1 92.37$\pm$1.0
5. Replace STPE-C with STPE-U 96.99$\pm$0.2 94.37$\pm$1.1 89.22$\pm$2.2 75.38$\pm$5.1 64.82$\pm$0.7 50.03$\pm$2.9 78.59$\pm$2.2 64.32$\pm$4.6
6. Remove mask in Eq. (11) 98.14$\pm$0.0 98.69$\pm$0.0 98.08$\pm$0.1 98.19$\pm$0.1 88.95$\pm$0.1 85.81$\pm$1.4 90.31$\pm$0.1 92.05$\pm$0.3
7. Use recent sampling 97.49$\pm$0.1 97.86$\pm$0.1 96.44$\pm$0.3 96.93$\pm$0.2 87.23$\pm$0.8 85.91$\pm$0.1 86.25$\pm$1.6 87.15$\pm$1.5
@@ 6. Conclusions
This paper introduces the Dynamic Graph Transformer with Correlated Spatial-Temporal Positional Encoding (DGTC), a novel approach for representation learning on Continuous-Time Dynamic Graphs (CTDGs). We employ the Poisson Point Process assumption and sampled temporal neighborhood to achieve comprehensive proximity estimation on CTDGs. Subsequently, we propose Correlated Spatial-Temporal Positional Encodings (STPE-C), which utilizes the comprehensive proximity to capture spatial-temporal high-order proximity. Extensive experiments conducted on seven small and two large-scale datasets demonstrate the performance superiority and scalability of the proposed DGTC model. A potential future direction for this work could involve designing more sophisticated spatial-temporal distances for improved proximity estimation and preservation.
@@ Appendix B More discussion on other Stochastic process
Appendix C Experimental Setting
C.1. Datasets
@@ C.2. Baselines
C.3. Hyper-parameters Tuning
@@
Improve 1.24 0.52 4.34 2.56 5.40 1.56 1.48
C.4. Implementation details of CorDGT
The code of CorDGT well be public available once this paper is accepted.
For fair comparison, we train all the models for 50 epochs with early stopping executed if there is no improvement on validation AP for 3 epochs. In addition, we use the batch size of 100 for all models. We repeat the methods for 3 runs and report the mean and standard deviation of statistics. For the proposed CorDGT, we set the learning rate as 0.001 and optimizer as Adam for all datasets. The attention heads, layer number and hidden dimension of main CorDGT encoder is set as 6, 2, 64, respectively, for all datasets. We sample 2-hop temporal neighbors with sampling number {20,1} for all datasets. We set the dimension of Correlated Spatial-Temporal Positional Encoding as 200, where the dimension of Spatial Distance encoding and Temporal distance is set as 100 and 100, respectively. We set $\alpha=1$ and $\beta=10$ for Reddit, Wikipedia and UCI, and set $\alpha=1,$ and $\beta=0.1$ for LastFM. All the experiments are run on a Linux Ubuntu 18.04 Server with a NVIDIA RTX2080Ti GPU.
Appendix D Additional Experimental results
D.1. Node classification results
@@
D.2. AUC results of link prediction
@@
D.3. Scalability on Gowalla Outdoors

### v2

@@ Abstract.
Learning effective representations for Continuous-Time Dynamic Graphs (CTDGs) has garnered significant research interest, largely due to its powerful capabilities in modeling complex interactions between nodes. A fundamental and crucial requirement for representation learning in CTDGs is the appropriate estimation and preservation of proximity. However, due to the sparse and evolving characteristics of CTDGs, the spatial-temporal properties inherent in high-order proximity remain largely unexplored. Despite its importance, this property presents significant challenges due to the computationally intensive nature of personalized interaction intensity estimation and the dynamic attributes of CTDGs. To this end, we propose a novel Correlated Spatial-Temporal Positional encoding that incorporates a parameter-free personalized interaction intensity estimation under the weak assumption of the Poisson Point Process. Building on this, we introduce the Dynamic Graph Transformer with Correlated Spatial-Temporal Positional Encoding (CorDGT), which efficiently retains the evolving spatial-temporal high-order proximity for effective node representation learning in CTDGs. Extensive experiments on seven small and two large-scale datasets demonstrate the superior performance and scalability of the proposed CorDGT. The code is available at: https://github.com/wangz3066/CorDGT.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanybooktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanydoi: 10.1145/3701551.3703489isbn: 979-8-4007-1329-3/25/03ccs: Theory of computation Dynamic graph algorithms
@@ 1. Introduction
Graph Neural Networks (GNNs) (Veličković et al., 2018; Hamilton et al., 2017; Xu et al., 2019; Dong et al., 2021; Wang et al., 2024b) have become a potent tool for analyzing diverse graph structures due to their effectiveness in learning low-dimensional graph representations. While early GNN research focused on static graphs, many real-world network data exhibit evolving graph structures, such as the World Wide Web and recommendation systems (Chen et al., 2023; Wang et al., 2024a; Wu et al., 2022). The increasing prevalence of dynamic graph data has spurred researchers to adapt GNNs to handle dynamic graphs. Dynamic graphs can be categorized into Discrete-Time Dynamic Graphs (DTDGs) and Continuous-Time Dynamic Graphs (CTDGs). Recently, CTDGs have garnered more research attention compared to DTDGs, attributed to their flexible generalization capabilities and proficiency in modeling intricate and dynamic node interactions. To learn effective representations for CTDGs, researchers have explored various techniques, including temporal message passing (Xu et al., 2020b; Rossi et al., 2020; Kumar et al., 2019; Trivedi et al., 2019; Wang et al., 2024c) and the incorporation of temporal intervals into random walks (Nguyen et al., 2018; Wang et al., 2021b; Jin et al., 2022), etc.
@@ 5.3. Ablation Studies
1. Set $\alpha=0$ in Eq. (7) 98.09$\pm$0.0 98.24$\pm$0.1 98.17$\pm$0.0 98.28$\pm$0.3 89.10$\pm$0.6 92.07$\pm$0.4 91.30$\pm$0.3 91.52$\pm$0.3
2. Set $\beta=0$ in Eq. (7) 98.10$\pm$0.0 98.19$\pm$0.1 97.87$\pm$0.0 98.18$\pm$0.2 89.55$\pm$0.6 90.80$\pm$1.7 90.46$\pm$0.3 94.28$\pm$0.8
3. Remove temporal distance 97.91$\pm$0.1 98.33$\pm$0.3 97.81$\pm$0.1 97.11$\pm$0.1 88.43$\pm$0.1 86.68$\pm$0.4 92.19$\pm$0.1 91.20$\pm$0.3
4. Remove spatial distance 97.81$\pm$0.0 98.02$\pm$0.1 97.62$\pm$0.0 97.36$\pm$0.4 92.47$\pm$0.1 91.43$\pm$0.5 90.15$\pm$0.1 92.37$\pm$1.0
5. Replace STPE-C with STPE-U 96.99$\pm$0.2 93.87$\pm$1.1 89.22$\pm$2.2 75.38$\pm$5.1 64.82$\pm$0.7 50.03$\pm$2.9 78.59$\pm$2.2 64.32$\pm$4.6
6. Remove mask in Eq. (11) 98.14$\pm$0.0 98.19$\pm$0.0 98.08$\pm$0.1 98.19$\pm$0.1 88.95$\pm$0.1 85.81$\pm$1.4 90.31$\pm$0.1 92.05$\pm$0.3
7. Use recent sampling 97.49$\pm$0.1 97.36$\pm$0.1 96.44$\pm$0.3 96.93$\pm$0.2 87.23$\pm$0.8 85.91$\pm$0.1 86.25$\pm$1.6 87.15$\pm$1.5
@@ 6. Conclusions
This paper introduces the Dynamic Graph Transformer with Correlated Spatial-Temporal Positional Encoding (CorDGT), a novel approach for representation learning on Continuous-Time Dynamic Graphs (CTDGs). We employ the Poisson Point Process assumption and sampled temporal neighborhood to achieve comprehensive proximity estimation on CTDGs. Subsequently, we propose Correlated Spatial-Temporal Positional Encodings (STPE-C), which utilizes the comprehensive proximity to capture spatial-temporal high-order proximity. Extensive experiments conducted on seven small and two large-scale datasets demonstrate the performance superiority and scalability of the proposed CorDGT model. A potential future direction for this work could involve designing more sophisticated spatial-temporal distances for improved proximity estimation and preservation.
ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China (62476244), Zhejiang Provincial Natural Science Foundation of China (Grant No: LTGG23F030005), National Natural Science Foundation of China (62372399,62476245) and the advanced computing resources provided by the Supercomputing Center of Hangzhou City University.
Ethical Considerations
The proposed CorDGT is used to learn the temporal embeddings which can be leveraged for downstream tasks such as link prediction and node classification. The direct negative societal effects of this research, encompassing fairness, privacy, and security considerations, are minimal. Nevertheless, akin to other predictive models, a few erroneous predictions by the model could impact system functionality. Despite extensive experimental validation of the model’s efficacy, occasional inaccurate predictions, particularly on outlier data, remain plausible. Therefore, enhancing data quality through measures such as data cleaning prior to model application is advised.
@@ Appendix B More discussion on other Stochastic process
Appendix C Complexity Comparison with existing models
In Section 4, we analyze the computation complexity of CorDGT . In this section, we also analyze the computation complexity of two baseline methods, i.e., TGAT (Xu et al., 2020b) and TGN (Rossi et al., 2020). For these models, given a mini-batch of interactions of size $B$, the historical neighbor sampling process costs $O(B\log(\bar{d}))$ , where $\bar{d}$ represents the average degree of nodes. The forward process, costs $O(CHD)$ where $C,H,D$ denote the number of sampled neighbors, attention heads, and hidden dimensions, respectively. Thus, the overall time complexity of TGAT and TGN is $(Blog(\bar{d})+CHD)$. The size of network parameters is denoted as $P$. In addition, TGN requires a memory and costs the spatial complexity of $O(ND^{\prime})$ where $N$ and $D^{\prime}$ denote the number of nodes and hidden dimension of the memory. We summarize the time and space complexity in Table 3.
Table 3. The time and space complexity of different models.
Space Complexity Time Complexity
TGAT $O(B\log(\bar{d})+CHD)$ $O(P)$
TGN $O(B\log(\bar{d})+CHD)$ $O(P+ND^{\prime})$
CorDGT $O(B(\log(\bar{d})+C)+C^{2}HD)$ $O(P+|\mathcal{E}|)$
Appendix D Experimental Setting
D.1. Datasets
@@ D.2. Baselines
D.3. Hyper-parameters Tuning
@@
Improve 1.24 0.52 4.34 1.06 5.40 1.56 1.48
D.4. Implementation details of CorDGT
For fair comparison, we train all the models for 50 epochs with early stopping executed if there is no improvement on validation AP for 3 epochs. In addition, we use the batch size of 100 for all models. We repeat the methods for 3 runs and report the mean and standard deviation of statistics. For the proposed CorDGT, we set the learning rate as 0.001 and optimizer as Adam for all datasets. The attention heads, layer number and hidden dimension of main CorDGT encoder is set as 6, 2, 64, respectively, for all datasets. We sample 2-hop temporal neighbors for all datasets. For LastFM and UCI, the sampling numbers of contextual neighbors are {32,1}, and for other datasets, the sampling numbers of contextual neighbors are {20,1}. We set $\alpha=0.1$ and $\beta=1.0$ for UCI and $\alpha=1$ and $\beta=0.1$ for LastFM. We set $\alpha=1$ and $\beta=10$ for other datasets. We set the dimension of Correlated Spatial-Temporal Positional Encoding as 200, where the dimension of Spatial Distance encoding and Temporal distance is set as 100 and 100, respectively. All the experiments are run on a Linux Ubuntu 18.04 Server with a NVIDIA RTX2080Ti GPU.
Appendix E Additional Experimental results
E.1. Node classification results
@@
E.2. AUC results of link prediction
@@
E.3. Scalability on Gowalla Outdoors